KTeV Shift Summary 11/20/99

Tony Barker (tonyb@cuhep.Colorado.EDU)
Sat, 20 Nov 1999 14:39:06 -0700 (MST)

Shift Summary for Day Shift, 11/20/99

On Shift: Hogan Nguyen, Tony Barker

Hours of beam available at all intensities: 8:15
Hours of beam available at nominal intensity: 8:15
Hours of E799 data (special runs not included): 8:15
Hours of special runs (laser scans, muon runs...): 0:00

Accounting of lost time: Switch to three turns from two
to raise intensity to 1.0E13.

========================================================================

Runs ended this shift: none (HOWEVER 15020 ran throughout shift,

15020 ended at 15:35 with 19.0 million events;
409k online 3pi0-Dalitz)

Tapes brought to FCC this shift: UPA157-159
UPB157-159
UPC157-159
UPS053

========================================================================

Accesses during shift: none

========================================================================

Overview of shift: Generally smooth E799 running.

Intensity was at first rather low (8.5E12).
BobT requests higher intensity; gets operators
to agree to change accelerator configuration
from two to three 'turns' if we agree to take
1.0E13 (this to avoid dumping too much beam).
Intensity request therefor changed to 1.0E13
at about 10:30.

Spills at 1.1E13 will cause DAQ_DEAD problems,
but 1.0E13 seems ok, with typical livetime
of approximately 64%.

Twenty-six more UPH Hyperon split tapes were
labelled.

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------
15202 156 07:00 Smooooth running at shift change.

314 10:30 Request intensity increase to 1.0E13.

341 11:05 Stable running at 1.0E13.

538 15:30 Shift ends, still smoooth running.

========================================================================
Items to Follow up on Next Shift

Call Sydney when run 15020 ends, BEFORE unloading the
tapes, so that he can check the amount of tape usage.

Look into VV-prime latch problems if there is any
down time. [e.g., problems seem to temporarily
disappear following KFREND; do they reappear when
beam is off?]